# Sense4FL: Vehicular Crowdsensing Enhanced Federated Learning for Autonomous Driving

### v1

@@ Abstract
To accommodate constantly changing road conditions, real-time model training is essential for autonomous driving (AD). Federated learning (FL) serves as a promising paradigm to enable autonomous vehicles to train models collaboratively with their onboard computing resources. However, existing vehicle selection schemes for FL all assume predetermined and location-independent vehicles’ datasets, neglecting the fact that vehicles collect training data along their routes, thereby resulting in suboptimal vehicle selection. To improve the perception quality in AD for a region, we propose Sense4FL, a vehicular crowdsensing-enhanced FL framework featuring trajectory-dependent vehicular training data collection. To this end, we first derive the convergence bound of FL by considering the impact of both vehicles’ uncertain trajectories and uploading probabilities, from which we discover that minimizing the training loss is equivalent to minimizing a weighted sum of local and global earth mover’s distance (EMD) between vehicles’ collected data distribution and global data distribution. Based on this observation, we formulate the trajectory-dependent vehicle selection and data collection problem for FL in AD. Given that the problem is NP-hard, we develop an efficient algorithm to find the solution with an approximation guarantee. Extensive simulation results have demonstrated the effectiveness of our approach in improving object detection performance compared with existing benchmarks.
@@ I Introduction
Autonomous driving (AD) enhances road safety, reduces traffic congestion, and provides environmental benefits, which has gained substantial attention lately [1, 2]. With joint efforts from the academia and auto industry, significant strides have been made in AD. For instance, multiple manufacturers, including Honda and Mercedes-Benz, start selling level 3 cars; Waymo offers rides in self-driving taxis to the public in Arizona (Phoenix) and California (San Francisco and Los Angeles) as of 2024 [3, 4]. Without any doubt, the aforementioned thrilling progress would not have been possible without the advancement of artificial intelligence (AI) capabilities on vehicles, which enable vehicles to detect objects, perceive their environment, and navigate in diverse scenarios.
@@
Despite significant progress in AD, its commercial use is still hampered by real-world deployment challenges and accidents. For instance, several Cruise vehicles were entangled in Muni wires and caution tapes since they failed to detect these objects in extreme weather [5]. A Cruise robotaxi struck a pedestrian and dragged her 20 feet as it failed to classify and track the pedestrian [6]. One primary reason for these accidents is that vision models, such as object classification and detection models, are known to lack generalization capabilities under changing environmental conditions and domain shifts, such as diverse street scenes and extreme weather conditions [7][8][9]. To enhance safety of AD, it is essential to adapt (fine-tune) a vision model for a specific region and improve it over time.
@@
In this paper, we explore a scenario where vehicles collectively collect street views and train (fine-tune) a vision model, e.g., object detection/classification model, for a region of interest, like a city. The trained (fine-tuned) model can then be distributed to all vehicles passing through this region to improve their perception accuracy. To achieve this, federated learning (FL) is a promising approach to training a global model collectively from a number of vehicles [10], where vehicles update and upload their local models to an FL server for aggregation. Compared with centralized learning, FL has the following salient advantages in the context of AD. First of all, FL may significantly reduce data upload volume. The raw data rate of a 1080p video stream from a typical vehicle camera can amount to 1493 Mbps [^1] We calculate the data rate by considering a color depth of 24 bits and a frame rate of 30 fps. [11], and each vehicle may be equipped with six or more cameras and other sensors, such as LiDAR. In comparison, the state-of-the-art object detection model Yolov8m has 25.9 million parameters [12], i.e., around 52 MB in 16-bit, which is considerably smaller than the sensory data size generated over a period of interest. Second, by leveraging the onboard computing capabilities of a large number of vehicles in parallel, FL is more scalable than centralized learning as it eliminates the need for a powerful central server. Finally, FL safeguards the location privacy and driving behaviors of drivers by preventing application servers from directly accessing their precise locations [^2] While the FL server can infer that the vehicle is in this region, the precise location can be protected. and driving states.
Given the advantages of FL for AD, many research efforts have been made to design FL schemes for vehicles [13, 14, 15]. However, all existing works consider location-independent vehicles’ datasets, akin to conventional FL settings where clients have predetermined local datasets [16, 17]. In reality, vehicles collect sensory data, such as street view information, along their routes, making them training data collectors with datasets dependent on their routes. Vehicle selection schemes ignoring this fact will result in model bias and poor training performance in FL [18]. In addition to vehicle selection, the data collection strategy must also be carefully designed. This is crucial not only for ensuring the model accurately represents the region but also for increasing the likelihood of successful model uploading. By jointly considering data collection and vehicle selection, we can optimize the FL process and enhance the performance of AD systems in the considered region.
@@
Based on the aforementioned observations, in this paper, we introduce a trajectory-dependent vehicular training data collection framework, i.e., vehicular crowdsensing enhanced FL (Sense4FL), to facilitate FL for AD. In this paper, we specifically focus on the object classification/detection task in AD, where vehicles serve as data collectors to capture images and videos of road scenes to constitute their training datasets and act as FL clients to train a global model for a region of interest collaboratively. To minimize training loss, our framework optimizes vehicle selection and data collection decisions by taking two factors into account: 1) the distribution of collected datasets according to selected vehicles’ routes, and 2) uploading probabilities resulting from vehicles’ routes, computing capabilities, and communication capabilities. Prior works on vehicular FL only consider the second factor, which may cause model performance degradation since the training datasets from participating vehicles may not reflect the global data distribution in the region of interest. The key contributions of this paper are summarized below.
@@
  • First, we present the Sense4FL framework for training an object classification/detection model in AD, which determines vehicle selection and data collection in a region of interest by considering street data distribution and vehicular mobility.
@@
  • Second, we derive the convergence bound of FL by considering the impact of both vehicles’ uncertain trajectories and uploading probabilities. We discover that minimizing the training loss is equivalent to minimizing a weighted sum of local and global earth mover’s distance (EMD) between vehicles’ collected data distribution and global data distribution.
@@
  • Then, we formulate the joint vehicle selection and training data collection problem to minimize FL training loss. It turns out that this problem is a non-linear multiple-choice knapsack problem with a non-convex non-separable objective function with non-continuous variables. Given that the optimization problem is NP-hard and highly challenging, we develop an efficient algorithm to obtain the solution with an approximation guarantee.
@@
  • Finally, we demonstrate the effectiveness of our approach for the state-of-the-art object detection model, YOLO, with the nuImages dataset. The simulation results show that our algorithm significantly improves the performance of object detection compared to existing benchmarks and enables fast fine-tuning of models under changing environments.
@@ II Related Work
Due to data and resource heterogeneity in FL [10, 13, 14, 15, 16, 18], client selection is crucial for performance. Numerous research efforts have been conducted in this area. In [19], Cho et al. performed a convergence analysis of FL over biased client selection, demonstrating that selecting clients with higher local loss achieves faster convergence. To optimize FL in wireless networks, Nishio et al. proposed a client selection scheme aimed at maximizing the number of uploaded models to enhance learning performance [16]. Chen et al. optimized client selection and radio resource allocation by taking packet errors into account [20]. Considering bandwidth limitations, Huang et al. in [21] optimized client selection with a fairness guarantee based on Lyapunov optimization. Xu et al. in [22] formulated a stochastic optimization problem for joint client selection and bandwidth allocation under long-term client energy constraints. By considering heterogeneous client hardware constraints and data quality, Deng et al. developed an automated, quality-aware client selection framework for FL [23]. Zhu et al. introduced an asynchronous FL framework with adaptive client selection to minimize training latency while considering client availability and long-term fairness [24]. However, these client selection schemes do not account for user mobility and are therefore not well-suited for vehicular environments.
@@
As a special case of mobile users, vehicles can act as FL clients to collectively train a model for vehicular applications [25, 26, 27, 28]. Along this line, Ye et al. proposed a contract-theory-based vehicle selection approach that accounts for image quality and heterogeneous vehicle capabilities [29]. In [30], Pervej et al. presented a resource allocation and client selection framework and analyzed its learning performance under full and partial vehicle participation. Zhao et al. proposed Newt, an improved vehicle selection mechanism with feedback control by considering data and resource heterogeneity in dynamic environments [31]. To ensure timely completion of FL iterations within latency constraints, Zhao et al. in [32] maximized vehicle participation by accounting for dynamic wireless channels and heterogeneous computing capacities. In [33], Xiao et al. proposed a min-max optimization framework that selects vehicles based on image quality while minimizing the overall system cost in FL. Furthermore, Zheng et al. [34] introduced AutoFed, a heterogeneity-aware FL framework that leverages multimodal sensory data in autonomous vehicles and incorporates a vehicle selection mechanism based on model similarities to enhance training stability and convergence.
@@
Although existing research has investigated FL in vehicular networks, most of the aforementioned works assume vehicle training data is location-independent, similar to traditional FL frameworks. However, vehicles proactively collect data from their surroundings, resulting in data distributions dependent on their routes. In FL, the data distribution of clients plays an essential role in learning performance, as deviation from the desired distribution can introduce biases during model training, resulting in severe accuracy degradation [18, 35, 36, 37]. By observing this, we propose a unified framework for trajectory-dependent vehicular data collection and training for FL, thereby significantly improving model performance in AD.
@@ III The Sense4FL Framework
In this section, we elaborate on the Sense4FL framework, including the vehicular mobility model, the training data collection, the federated learning procedure, and the overview of Sense4FL.
Figure 1: Illustration of Sense4FL. Each autonomous vehicle acts as a mobile data collector and an FL client. As a vehicle traverses through a region, it collects data about street blocks and then leverages the collected data to train local models for FL. The selection of vehicles participating in the training process and the timing of starting their training are carefully designed by considering the impact of vehicles’ uncertain trajectories.
@@ III-A System Model
III-A1 The vehicular mobility model
As illustrated in Fig. 1, we consider a region of interest consisting of multiple street blocks, where an FL server is responsible for model aggregation. We assume vehicles can upload their models to the FL server via cellular networks, i.e., any associated base station (BS), which then forwards the models to the FL server for aggregation through wired links. Let =$\mathcal{B}=\{1,2,\dots,B\}$ and $\mathcal{V}=\{1,2,\dots,V\}$ denote the sets of street blocks and vehicles, respectively. The set of sojourn time for vehicles in each street block can be represented by $\mathcal{T}=\{t_{1,1},t_{1,2},...,t_{V,B}\}$, where $t_{v,b}$ denotes the sojourn time of vehicle $v$ in street block $b$, which follows a probability distribution, e.g., a truncated Gaussian distribution as often adopted in [38][39][40]. $t_{v,b}$ can be estimated by the FL server based on historical traffic information. Each autonomous vehicle serves as both a mobile data collector and an FL client, which collects training data from road environments, i.e., images of road conditions, traffic signs, pedestrians, and vehicles, upon traversing through a street block. The FL procedure selects a subset of vehicles from this region, leveraging their onboard computing capabilities and locally collected datasets to train an object classification/detection model to enhance AD performance.
As shown in Fig. 1, a vehicle’s trajectory can be modeled as a sequence of street blocks. The trajectory of vehicle $v$ can be denoted by $h_{v}=\{s_{v}^{1},s_{v}^{2},...,s_{v}^{N_{v}}\}$, where $s_{v}^{n}$ represents the $n$-th element in this route with $N_{v}$ street blocks in total. Each $s_{v}^{n}$ corresponds to a street block in the set $\mathcal{B}$. In practice, an FL server cannot directly know the exact itinerary of a vehicle but can only predict the vehicle’s trajectories based on its location and driving direction. Consequently, a vehicle may have multiple potential routes. Specifically, the set of possible trajectories for vehicle $v$ can be expressed as $\mathcal{H}_{v}=\{h_{v,1},h_{v,2},\ldots,h_{v,M_{v}}\}$, where $h_{v,m}=\{s_{v,m}^{1},s_{v,m}^{2},\ldots,s_{v,m}^{N_{v,m}}\}$ represents the $m$-th potential trajectory out of $M_{v}$ trajectories. The probability of vehicle $v$ to choose trajectory $h_{v,m}$ is $q_{v,m}$.
@@ III-A2 Training Data Collection
We consider a multi-class object classification/detection task [41, 42]. Each training data sample is represented by $(\mathbf{x},y)$. Here, $\mathbf{x}\in\mathbb{R}^{d}$ is the input feature vector belonging to a compact space $\mathcal{X}$ whereas $y\in\mathbb{R}$ is the corresponding label from the label set $\mathcal{Y}=\{1,2,\ldots,C\}$. With different street scenes, we assume different street blocks follow heterogeneous data distribution, e.g., with different numbers of pedestrians, vehicles, or other objects. By defining the probability of a data sample in street block $b$ belonging to class $i$ as $p_{b}^{i}$, the global data distribution for the entire region can be expressed as follows
@@
where ${Q}_{b}$ is the average number of objects a vehicle encounters in street block $b$. It is noted that $p_{b}^{i}$ can be estimated in our systems based on public-domain information, such as satellite imagery and traffic information, or by requesting some vehicles to upload the statistical data (instead of the raw data) of the street block.

### v2

# Sense4FL: Vehicular Crowdsensing Enhanced Federated Learning for Object Detection in Autonomous Driving

@@ Abstract
To accommodate constantly changing road conditions, real-time vision model training is essential for autonomous driving (AD). Federated learning (FL) serves as a promising paradigm to enable autonomous vehicles to train models collaboratively with their onboard computing resources. However, existing vehicle selection schemes for FL all assume predetermined and location-independent vehicles’ datasets, neglecting the fact that vehicles collect training data along their routes, thereby resulting in suboptimal vehicle selection. In this paper, we focus on the fundamental perception problem and propose Sense4FL, a vehicular crowdsensing-enhanced FL framework featuring trajectory-dependent vehicular training data collection to improve the object detection quality in AD for a region. To this end, we first derive the convergence bound of FL by considering the impact of both vehicles’ uncertain trajectories and uploading probabilities, from which we discover that minimizing the training loss is equivalent to minimizing a weighted sum of local and global earth mover’s distance (EMD) between vehicles’ collected data distribution and global data distribution. Based on this observation, we formulate the trajectory-dependent vehicle selection and data collection problem for FL in AD. Given that the problem is NP-hard, we develop an efficient algorithm to find the solution with an approximation guarantee. Extensive simulation results have demonstrated the effectiveness of our approach in improving object detection performance compared with existing benchmarks.
@@ I Introduction
Autonomous driving (AD) enhances road safety, reduces traffic congestion, and provides environmental benefits, which has gained substantial attention lately [1, 2, 3]. With joint efforts from the academia and auto industry, significant strides have been made in AD. For instance, multiple manufacturers, including Honda and Mercedes-Benz, start selling level 3 cars; Waymo offers rides in self-driving taxis to the public in Arizona (Phoenix) and California (San Francisco and Los Angeles) as of 2024 [4, 5]. Despite significant progress in AD, its commercial use is still hampered by real-world deployment challenges and accidents. For instance, several Cruise vehicles were entangled in Muni wires and caution tapes since they failed to detect these objects in extreme weather [6]. A Cruise robotaxi struck a pedestrian and dragged her 20 feet as it failed to classify and track the pedestrian [7]. One primary reason for these accidents is that vision models, such as object classification and detection models, are known to lack generalization capabilities under changing environmental conditions and domain shifts, such as diverse street scenes and extreme weather conditions [8]. To enhance the safety of AD, it is essential to adapt a vision model for a specific region and improve it over time.
@@
To improve models on the fly, federated learning (FL) [9] serves as a promising paradigm, where vehicles update and upload their local models to a server for aggregation. Compared with centralized learning, FL has the following salient advantages in the context of AD. First of all, FL may significantly reduce data upload volume. The raw data rate of a 1080p video stream from a typical vehicle camera can amount to 1493 Mbps [^1] We calculate the data rate by considering a color depth of 24 bits and a frame rate of 30 fps. [10], and each vehicle may be equipped with six or more cameras and other sensors, such as LiDAR. In comparison, the state-of-the-art object detection model Yolov8m has 25.9 million parameters [11], i.e., around 52 MB in 16-bit, which is considerably smaller than the sensory data size generated over a period of interest. Second, by leveraging the onboard computing capabilities of a large number of vehicles in parallel, FL is more scalable than centralized learning as it eliminates the need for a powerful central server. Finally, FL safeguards the location privacy and driving behaviors of drivers by preventing application servers from directly accessing their precise locations [^2] While the FL server can infer that the vehicle is in this region, the precise location can be protected. and driving states.
Given the advantages of FL for AD, many research efforts have been made to design FL schemes for vehicles [12, 13]. However, all existing works consider location-independent vehicles’ datasets, akin to conventional FL settings where clients have predetermined local datasets [14, 15]. Nevertheless, considering a FL scenario where vehicles collectively collect street views and train (adapt) a vision model, e.g., object detection/classification model, for a region of interest, like a city [^3] Our proposed framework can be extended to other critical tasks, such as semantic segmentation and control command (acceleration/deceleration, go/stop), which can be investigated in future work., vehicles collect sensory data, such as street view information, along their routes. Vehicle selection schemes without explicitly considering their trajectory-dependent data distributions result in inferior performance in FL. To illustrate the effect of the trajectory-aware vehicle selection, we consider a simple case with four street blocks forming three trajectories, represented by $h_{1}=\{1\}$, $h_{2}=\{2,4,3\}$, $h_{3}=\{2,4\}$. Some street blocks may have more cars whereas some may have more pedestrians, as shown in Fig. 1(a). We evaluate FL performance under several vehicle selection strategies based on trajectories, assuming that each vehicle follows one of these three trajectories. We also assume sufficient communication-computing resources so that trajectories only affect the data distributions without impacting other aspects, e.g., the model uploading probability. As shown in Fig. 1(b), appropriate trajectory-aware vehicle selection can outperform trajectory-agnostic random selection by $4.4\%$, because it can select vehicles with more representative data for this region.
(a) The normalized frequency of
objects for each street block.
(b) Final test accuracy for different selection strategies.
Figure 1: The data distribution of four neighboring street blocks and the final test accuracy (after convergence) via different vehicle selection strategies in the nuImages dataset. $\{h_{x},h_{y}\}$ means that we select one vehicle from trajectory $x$ and one vehicle from $y$ in each round; “Random” means we randomly choose two vehicles in each round.
@@
Based on the aforementioned observations, in this paper, we introduce a trajectory-dependent vehicular training data collection framework, i.e., vehicular crowdsensing enhanced FL (Sense4FL), to facilitate FL for AD. We first rigorously derive the convergence upper bound of Sense4FL by considering the trajectories of selected vehicles, characterizing the upper bound using earth mover’s distance (EMD) metrics in terms of vehicles’ and global data distributions. To minimize the upper bound, our framework optimizes vehicle selection and data collection decisions by taking two factors into account: i) the distribution of collected datasets according to selected vehicles’ routes, and ii) uploading probabilities resulting from vehicles’ routes, computing capabilities, and communication capabilities. Note that prior works on vehicular FL neglect the first aspect, which may cause model performance degradation since the training datasets from participating vehicles may not reflect the global data distribution in the region of interest. The key contributions of this paper are summarized below.
@@
  • First, we present the Sense4FL framework for training an object detection model in AD, which determines vehicle selection and data collection in a region of interest by considering street data distribution and vehicular mobility.
@@
  • Second, we derive the convergence bound of FL by considering the impact of both vehicles’ uncertain trajectories and uploading probabilities, establishing the theoretical relationship between model accuracy loss and data collection. We discover that minimizing the training loss is equivalent to minimizing a weighted sum of client and global EMD between vehicles’ collected data distribution and regional data distribution.
@@
  • Then, we formulate the joint vehicle selection and training data collection problem to minimize FL training loss. It turns out that this problem is a non-linear multiple-choice knapsack problem (MCKP) with a non-convex non-separable objective function with non-continuous variables. Given that the optimization problem is NP-hard and highly challenging, we develop an efficient algorithm to obtain the solution with an approximation guarantee.
@@
  • Finally, we demonstrate the effectiveness of our approach for the state-of-the-art object detection model, YOLO, with the nuImages dataset. The simulation results show that our algorithm significantly improves the performance of object detection compared to existing benchmarks and enables fast adaptation of models under changing environments.
@@ II Related Work
A substantial body of research has focused on optimizing federated learning at the network edge [16, 17, 18, 19]. Due to data and resource heterogeneity in FL [9, 12, 14], client selection is crucial for learning performance. Numerous research efforts have been conducted in this area. Nishio et al. in [14] proposed a client selection scheme in wireless networks aimed at maximizing the number of uploaded models to enhance learning performance. In [20], Cho et al. performed a convergence analysis of FL over biased client selection, demonstrating that selecting clients with higher local loss achieves faster convergence. Chen et al. optimized client selection and radio resource allocation by taking packet errors into account [21]. Considering bandwidth limitations, Huang et al. in [22] optimized client selection with a fairness guarantee based on Lyapunov optimization. Xu et al. in [23] formulated a stochastic optimization problem for joint client selection and bandwidth allocation under long-term client energy constraints. By considering heterogeneous client hardware constraints and data quality, Deng et al. developed an automated, quality-aware client selection framework for FL [24]. Zhu et al. introduced an asynchronous FL framework with adaptive client selection to minimize training latency while considering client availability and long-term fairness [25]. By identifying and excluding adverse local updates, Wu et al. proposed an optimal aggregation algorithm and a probabilistic client selection framework to accelerate model convergence [26]. However, these client selection schemes do not account for user mobility that may affect FL performance, which are not suited for vehicular environments.
TABLE I: Summary of related works in vehicular FL systems.
Ref.
Client
Selection
Data
Collection
Theoretical
Analysis
AD
Dataset
[27]
[28]
[29]
[30]
[31]
[32]
[33]
[34]
Ours
@@
As a special case of mobile users, vehicles can act as FL clients to collectively train a model for vehicular applications. Along this line, Ye et al. in [27] proposed a contract-theory-based vehicle selection approach that accounts for image quality and heterogeneous vehicle capabilities. In [28], Pervej et al. presented a resource allocation and client selection framework and analyzed its learning performance under full and partial vehicle participation. Zhao et al. proposed Newt, an improved vehicle selection mechanism with feedback control by considering data and resource heterogeneity in dynamic environments [29]. To ensure timely completion of FL iterations within latency constraints, Zhao et al. in [30] maximized vehicle participation by accounting for dynamic wireless channels and heterogeneous computing capacities. In [31], Xiao et al. proposed a min-max optimization framework that selects vehicles based on image quality while minimizing the overall system cost in FL. Zhang et al. in [32] introduced a mobility- and channel dynamic-aware FL scheme, which enables road side unit (RSU) to select appropriate vehicles and weightedly average the local models to improve the FL performance in vehicular networks. Zhang et al. in [33] investigate the joint optimization of vehicle selection, training time, and model quantization of FL with gradient quantization in vehicle edge computing by considering the mobility and the uncertainty of channel conditions. Furthermore, Zheng et al. [34] introduced AutoFed, a heterogeneity-aware FL framework that leverages multimodal sensory data to improve object detection performance in autonomous vehicles and incorporates a vehicle selection mechanism based on model similarities to enhance training stability.
@@
However, the aforementioned works assume vehicle training data is location-independent, similar to traditional FL frameworks. In reality, since vehicles proactively collect data from their surroundings, their data distributions are dependent on their routes. In FL, the data distribution of clients plays an essential role in learning performance, as deviation from the desired distribution can introduce biases during model training, resulting in severe accuracy degradation [35, 36]. To fill this research gap, this work provides a rigorous convergence analysis and devises a unified framework for trajectory-aware vehicle selection and training data collection to enhance learning performance in vehicular FL. To compare our work and related works, we provide a summary table in Table I.
@@ III The Sense4FL Framework
In this section, we elaborate on the Sense4FL framework, including the system model, specifically the vehicular mobility and the training data collection model, the federated learning procedure, and the overview of Sense4FL.
Figure 2: Illustration of Sense4FL framework. Each autonomous vehicle acts as a mobile data collector and an FL client. As a vehicle traverses through a region, it collects data about street blocks and then leverages the collected data to train local models for FL. The selection of vehicles participating in the training process and the timing of starting their training are carefully designed by considering the impact of vehicles’ uncertain trajectories.
@@ III-A System Model
As illustrated in Fig. 2, we consider a region of interest consisting of multiple street blocks, where an FL server is responsible for model aggregation. Each autonomous vehicle serves as both a mobile data collector and an FL client, which collects training data from road environments, i.e., images of road conditions, traffic signs, pedestrians, and vehicles, upon traversing through a street block. The FL procedure selects a subset of vehicles from this region, leveraging their onboard computing capabilities and locally collected datasets to train an object detection model to enhance AD performance. We assume vehicles can upload their models to the FL server via cellular networks, i.e., any associated base station (BS), which then forwards the models to the FL server for aggregation through wired links.
III-A1 Vehicular mobility
[…]
@@ III-A2 Training data collection
We consider a multi-class object detection task [39]. Each training data sample is represented by $(\mathbf{x},y)$. Here, $\mathbf{x}\in\mathbb{R}^{d}$ is the input feature vector belonging to a compact space $\mathcal{X}$ whereas $y\in\mathbb{R}$ is the corresponding label from the label set $\mathcal{Y}=\{1,2,\ldots,C\}$. With different street scenes, we assume different street blocks follow heterogeneous data distribution, e.g., with different numbers of pedestrians, vehicles, or other objects. By defining the probability of a data sample in street block $b$ belonging to class $i$ as $p_{b}^{i}$, the global data distribution for the entire region can be expressed as follows
@@
where ${Q}_{b}$ is the average number of objects a vehicle encounters in street block $b$. It is noted that $p_{b}^{i}$ can be estimated in our systems based on public-domain information, such as satellite imagery and traffic information, or by requesting some vehicles to upload the statistical data (instead of the raw data) of the street block [^5] In this paper, we consider data heterogeneity at the “street block” level. However, the proposed Sense4FL framework can also be applied to broader spatial scales, e.g., distinctions between downtown and residential areas..